OpenStudy (anonymous):

#40 is the sequence geometric ? if so identify the common ratio. 6, 12, 24, 48 A. yes 2 B. yes -2 C. yes -3 D. no

OpenStudy (anonymous):

Yes, it is A. 2, because you multiply by 2 every time. :)

OpenStudy (anonymous):

yes;2 6*2=12 12*2=24 24*2=48